OpenStudy (anonymous):

What is the value of 10-2 in standard form?

OpenStudy (anonymous):

1*10=10 2=2*10^0

OpenStudy (anonymous):

do you find the answer

OpenStudy (anonymous):

you must have a one digite number and multiple it by 10^m

OpenStudy (anonymous):

Do you mean 10^(-2) ? If that's what you are asking for, then that is equal to 1/10^2 = 1/100 = 0.01

OpenStudy (anonymous):

im confused whats the answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The question is 10 to the power of negative 2 in standard form

OpenStudy (anonymous):

whats the value

OpenStudy (anonymous):

With that clarification of yours, I stand by my original answer and steps.

OpenStudy (anonymous):

well can you explained to what standard form is because im not sure of that either

OpenStudy (anonymous):

\[10^{-2} = \frac{ 1 }{ 10^{2} } = \frac{ 1 }{ 100 } = 0.01\]They are all acceptable forms. There is no "standard" form, only simplifications.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! And you're welcome! @twitch